# MEGA-Bench[Uncaptioned image]: Scaling Multimodal Evaluation to over 500 Real-World Tasks

### v2

@@ 4 Experiments
We evaluate 22 VLMs with multi-image support on MEGA-Bench. §4.1 describes the models being assessed and the evaluation pipeline. §4.2 presents the quantitative evaluation results with a fine-grained analytical breakdown. §4.3 discusses the impact of the number of examples per task on the evaluation results. §4.4 provides error analysis based on GPT-4o. Detailed inspections of model behaviors with concrete error examples can be found in §F of the Appendix.
@@ 4.1 Evaluation Settings
Evaluated models. We evaluate a diverse range of large multimodal models. The proprietary models assessed include GPT-4o (0513) and GPT-4o mini (OpenAI, 2024a), Claude-3.5-Sonnet (0620 and 1022) (Anthropic, 2024a; b), Gemini-1.5-Pro (002) and Gemini-1.5-Flash (002) (Google, 2024a). For open-source models, we mainly focus on large flagship ($>$70B parameters) and small-to-medium efficiency models. The large models include Qwen2-VL-72B (Alibaba, 2024), InternVL2-Llama3-76B (Chen et al., 2024d), LLaVA-OneVision-72B (Li et al., 2024a), and NVLM (Dai et al., 2024). The medium-scale models comprise Qwen2-VL-(7B/2B) (Alibaba, 2024), Pixtral 12B (Mistral, 2024), Aria (Li et al., 2024d), InternVL2-(8B/2B) (Chen et al., 2024d), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V2.6 (Yao et al., 2024), LLaVA-OneVision-7B (Li et al., 2024a), Llama-3.2-11B Meta (2024), Idefics3-8B-Llama3 (Laurençon et al., 2024), and Aquila-VL-2B-llava-qwen (Gu et al., 2024).
Evaluation pipeline. MEGA-Bench has diverse and flexible formats. To ensure the models have clear instructions on the output format, we provide all evaluated VLMs with a one-shot in-context example. For each query, we fill in a pre-defined prompt template with the task instructions written by the task annotators, the 1-shot example, and the concrete query question. Since this one-shot example’s primary purpose is to illustrate the output format, we allocate only a tiny portion of the total image budget for it. For each model, we conduct experiments with and without Chain-of-Thought (CoT) prompting (Wei et al., 2022) for the Core tasks (the one-shot example of Open-ended tasks already contains CoT demonstrations). The prompt templates and other evaluation details (e.g., the frame sampling strategy for video inputs) are in §D.
Single-image setting. Our evaluation pipeline focuses on models with multi-image support. To properly evaluate models trained mainly for single-image use cases, we create a single-image setting using the single-image tasks of MEGA-Bench. See §A for the detailed results and analyses.
Table 2: The main results of different models on the Core and Open-ended subset of MEGA-Bench, with 440 and 65 tasks, respectively. We report the macro mean scores across all tasks in each set. The overall score is the weighted average of the Core and Open-ended scores. When computing the overall score, we use the higher Core score from ‘w/o CoT’ and ‘w/ CoT’. Some open-source models perform more reasonably when the input contains only a single image. Please see §A for results and analyses of our single-image setting.
@@
Aquila-VL-2B-llava-qwen (Gu et al., 2024) Efficiency Yes 16.32 16.00 24.57 17.38
InternVL2-2B (Chen et al., 2024d) Efficiency Yes 9.09 13.14 23.86 14.52
@@ 4.2 Main Results with breakdown analysis
Flagship models. Unlike the results on recent benchmarks like MMMU-Pro (Yue et al., 2024b) where GPT-4o (0513) and Claude-3.5-Sonnet (0620) get close scores, GPT-4o (0513) outperforms Claude-3.5-Sonnet (0620) with a clear margin on MEGA-Bench ($>$ 2%). Investigating the breakdown results, we observe that GPT-4o (0513) wins in most applications/skills except for coding, math, and planning-related tasks, where the answers are typically in a “structured” output format, as shown in Figure 5. The recent update of Claude-3.5-Sonnet (1022) makes consistent improvements across almost all dimensions, especially in planning-related tasks and those with infographics/UI/photographs inputs, and slightly surpasses GPT-4o (0513) in the overall score ($<$ 0.1%). The “Planning” application keyword contains tasks like symbolic planning (Zhu et al., 2024), navigation (Ku et al., 2020), chess games (Fu et al., 2024b), puzzle games (e.g., maze, Sudoku, etc.), and even the best models get low scores.
@@
Chain-of-Thought. An interesting finding is that the CoT prompt (See §D) effectively guides all proprietary models to generate a detailed reasoning process, and flagship-tier proprietary models all obtain better performance on the Core set. However, it has almost no effect on most open-source models. For example, the Qwen2-VL, InternVL2, and LLaVA-OneVision models rarely produce reasoning when given a CoT instruction, and sometimes get confused about the required format after generating the reasoning process, leading to a lower score on the Core set.
@@
Some open-source models get comparatively low scores for their parameter count. Llama-3.2-11B has difficulty leveraging the one-shot example to understand the correct output format and tends to generate a long descriptive sentence instead. This issue is alleviated under the CoT setting because the prompt provides extra instructions on the output format beyond the one-shot example, requesting the model to strictly separate the reasoning process from the final answer. Idefics3 frequently repeats the example answer from the one-shot demonstration. We suspect the reason for this problem is the poor support for multi-image (as our query contains at least two images, including the one-shot example) since it can generate reasonable responses with a single-image input. §A in the Appendix presents a single-image setting of MEGA-Bench and conducts further analyses.
4.3 Analysis on the number of samples per task
@@
As discussed in §1, one of MEGA-Bench’s goals is to optimize the inference cost while still producing detailed multi-dimensional breakdown analysis. Therefore, we prioritize expanding the number of tasks over adding many examples per task in the benchmark construction process. To understand the robustness of the benchmark score with around 15 examples per task, we obtained bootstrap distributions (Efron & Tibshirani, 1994; Hesterberg, 2011) of the model scores for our Core set with the CoT prompting. We did this by taking a random subset of the model’s responses of size $n$ ($n=3,5,\dots,13,15$) with replacement for each task and calculating the task-level macro-mean scores. To ensure the bootstrap distribution was numerically stable, we ran 10,000 Monte Carlo simulations. Figure 7 (left) shows that the variance in model scores rapidly narrows as the number of examples per task increases. As the number of examples per task increases beyond 7, the marginal return in variance reduction diminishes.
@@
Figure 7: (Left) The bootstrap distribution of benchmark scores of a subset of models as we gradually increase the bootstrap sample size of the number of examples per task. We use the results with the CoT prompting. Claude 3.5 Sonnet (1022) and Gemini 1.5 Pro (002) are used in this figure. (Right) The task-wise error distribution of GPT-4o (0513) over a subset of 255 Core tasks.
4.4 Error Analysis
@@
To understand the limitations of state-of-the-art VLMs, we analyze the GPT-4o (0513) results by manually identifying the error types over a subset of 255 tasks from the Core set. We use the CoT setting since the reasoning process helps determine the error type. Figure 7 (right) presents the error distribution. For GPT-4o, the lack of various reasoning capabilities (e.g., symbolic reasoning for planning/coding tasks, spatial or temporal reasoning for complex perception tasks, etc.) is the dominating failure mode on MEGA-Bench. Please refer to §F for the full definition of error types and detailed example-wise inspection results with different models.
@@ Appendix H Author Contribution Statement
All authors contributed at least 30 tasks to MEGA-Bench and participated in task brainstorming, annotation, and validation. They also engaged in discussions on data annotation and provided feedback on the project. The following authors made additional contributions to various aspects of the project:
Jiacheng Chen co-designed the project with Wenhu Chen, led the benchmark construction process, and coordinated collaboration among all contributors. Jiacheng Chen created and maintained the annotation GUI tool, GitHub repository, draft task taxonomy tree, and results visualization page to facilitate data annotation and improve data quality. Jiacheng Chen led the development and maintenance of the benchmark evaluation pipeline, including querying various VLMs and customizing evaluation metrics, and conducted the main experiments and analyses. Jiacheng Chen also led the writing of the paper, coordinating core contributors and incorporating their input into the manuscript.
Tianhao Liang co-led the benchmark data organization, implemented most of the model query pipelines under a consistent and unified framework, and conducted major experiments and analyses with Jiacheng Chen. Tianhao Liang helped maintain the evaluation pipeline and implemented the code execution metric. Tianhao Liang made significant efforts in data quality control, error analysis, and creating figures and tables for the paper.
Sherman Siu made major contributions to the benchmark construction process, including task reviewing, annotator coordination, data quality control, metric implementation, and code maintenance. Sherman Siu contributed and designed a bunch of complex and novel planning tasks. Sherman Siu also analyzed the number of examples per task to investigate the variance of the benchmark score and contributed to writing the main paper.
Zhengqing Wang contributed around 40 tasks, including several complex traditional computer vision tasks. Zhengqing Wang organized the benchmark statistics for the Appendix, participated in error case analysis, and developed the interactive functionalities of the project page.

### v3

@@ 4 Experiments
We evaluate 19 VLMs with multi-image support on MEGA-Bench. §4.1 describes the evaluated models and the evaluation pipeline. §4.2 presents the evaluation results with a fine-grained analytical breakdown. §4.3 provides analyses on the number of examples per task and error types.
@@ 4.1 Evaluation Settings
Evaluated models. We evaluate a diverse range of large multimodal models. The proprietary models assessed include GPT-4o (0513) and GPT-4o mini (OpenAI, 2024a), Claude-3.5-Sonnet (0620 and 1022) (Anthropic, 2024a; b), Gemini-1.5-Pro (002) and Gemini-1.5-Flash (002) (Google, 2024a). For open-source models, we mainly focus on large flagship ($>$70B parameters) and small-to-medium efficiency models. The large models include Qwen2-VL-72B (Alibaba, 2024), InternVL2-Llama3-76B (Chen et al., 2024d), LLaVA-OneVision-72B (Li et al., 2024a), and NVLM (Dai et al., 2024). The medium-scale models comprise Qwen2-VL-7B (Alibaba, 2024), Pixtral 12B (Mistral, 2024), Aria (Li et al., 2024d), InternVL2-8B (Chen et al., 2024d), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V2.6 (Yao et al., 2024), LLaVA-OneVision-7B (Li et al., 2024a), Llama-3.2-11B Meta (2024), and Idefics3-8B-Llama3 (Laurençon et al., 2024).
Evaluation pipeline. MEGA-Bench has diverse and flexible formats. To ensure the models have clear instructions on the output format, we provide all evaluated VLMs with a one-shot in-context example. For each query, we fill in a pre-defined prompt template with the task instructions written by the task annotators, the 1-shot example, and the concrete query question. Since this one-shot example’s primary purpose is to illustrate the output format, we allocate it a tiny portion of the total image budget. For each model, we conduct experiments with and without Chain-of-Thought (CoT) prompting (Wei et al., 2022) for the Core tasks. Full evaluation details are in §D. Our default evaluation pipeline focuses on models with multi-image support. To properly evaluate models trained mainly for single-image use cases, we create a single-image setting using the single-image tasks of MEGA-Bench. See §A for the detailed results and analyses of the single-image setting.
Table 2: The main results of different models on the Core and Open-ended subset of MEGA-Bench, with 440 and 65 tasks, respectively. We report the macro mean scores across all tasks in each set. The overall score is the weighted average of the Core and Open-ended scores. When computing the overall score, we use the higher Core score from ‘w/o CoT’ and ‘w/ CoT’.
@@ 4.2 Main Results with breakdown analysis
Flagship models. Unlike the results on recent benchmarks like MMMU-Pro (Yue et al., 2024b) where GPT-4o (0513) and Claude-3.5-Sonnet (0620) get close scores, GPT-4o (0513) outperforms Claude-3.5-Sonnet (0620) with a clear margin on MEGA-Bench ($>$ 2%). Investigating the breakdown results, we observe that GPT-4o (0513) wins in most applications/skills except for coding, math, and planning-related tasks, where the answers are typically in a “structured” output format ( Figure 5). The recent update of Claude-3.5-Sonnet (1022) makes improvements across almost all dimensions, especially in planning tasks and those with infographics/UI/photographs inputs, and slightly surpasses GPT-4o in the overall score ($<$ 0.1%). The “planning” application keyword contains tasks like symbolic planning (Zhu et al., 2024), navigation (Ku et al., 2020), chess games (Fu et al., 2024b), puzzle games (e.g., maze, Sudoku), etc., and even the best models get low scores.
@@
Chain-of-Thought. An interesting finding is that the CoT prompt (See §D) effectively guides all proprietary models to generate a detailed reasoning process, and flagship-tier proprietary models all obtain better performance on the Core set. However, it has almost no effect on most open-source models. For example, the Qwen2-VL, InternVL2, and LLaVA-OneVision models rarely produce reasoning with a CoT instruction, and sometimes get confused about the required format after generating the reasoning process, leading to a lower score on the Core set.
@@
Some open-source models get comparatively low scores for their parameter count (e.g., Llama-3.2-11B, Idefics3). One typical difficulty is leveraging the one-shot example to understand the output format, which is alleviated with CoT because the prompt provides extra instructions on the output format beyond the one-shot example. The lack of multi-image ability is another bottleneck of these models, and §A presents a single-image setting of MEGA-Bench with further analyses.
4.3 More Analysis
@@
Number of samples per task. As discussed in §1, one of MEGA-Bench’s goals is optimizing the inference cost while producing comprehensive breakdown analyses, making us prioritize the task diversity over examples per task in the benchmark construction process. To understand the robustness of the benchmark score with 15 examples per task, we obtained bootstrap distributions (Efron & Tibshirani, 1994; Hesterberg, 2011) of the model scores for our Core set with CoT. We did this by taking a random subset of the model’s responses of size $n$ ($n=3,5,\dots,15$) with replacement for each task and calculating the task-level macro-mean scores. To ensure the bootstrap distribution was numerically stable, we ran 10,000 Monte Carlo simulations. Figure 7 (left) shows that the variance in model scores rapidly narrows as the number of examples per task increases. As the number of examples per task increases beyond 7, the marginal return in variance reduction diminishes.
@@
Figure 7: (Left) The bootstrap distribution of benchmark scores (w/ CoT prompting) with a gradually increased bootstrap sample size of the number of per-task examples. (Right) The task-wise error distribution of GPT-4o (0513) over a subset of 255 Core tasks.
@@
Error analysis. To understand the limitations of state-of-the-art VLMs, we analyze the GPT-4o (0513) results by manually identifying the error types over a subset of 255 tasks from the Core set. We use the CoT setting since the reasoning process helps determine the error type. Figure 7 (right) presents the error distribution. For GPT-4o, the lack of various reasoning capabilities (e.g., symbolic reasoning for planning/coding tasks, spatial or temporal reasoning for complex perception tasks, etc.) is the dominating failure mode on MEGA-Bench. Please refer to §F for the full definition of error types and detailed example-wise inspection results with different models.
@@ 5 Conclusion
Acknowledgment
We would like to thank Vector Institute for their support throughout this research. We also express our gratitude to Tung Vu from GreenNode for providing access to GPUs, which were instrumental in running some of the evaluation experiments. Additionally, we appreciate the contributions of all open-source language model providers, whose efforts have significantly propelled the advancement of research in this field.
@@ Appendix H Author Contribution Statement
All authors contributed at least 30 tasks to MEGA-Bench and participated in task brainstorming, annotation, and validation. They also engaged in discussions on data annotation and provided feedback. The following authors made additional contributions to various aspects of the project:
Jiacheng Chen co-designed the project with Wenhu Chen, led the benchmark construction process, and coordinated collaboration among all contributors. Jiacheng Chen created and maintained the annotation GUI tool, GitHub repository, draft task taxonomy tree, and results visualization page to facilitate data annotation and improve data quality. Jiacheng Chen led the development and maintenance of the benchmark evaluation pipeline, including the VLM query pipeline, customized evaluation metrics, HuggingFace Space for the project, etc., and conducted the main experiments and analyses. Jiacheng Chen also led the writing of the paper, coordinating core contributors and incorporating their input into the manuscript.
Tianhao Liang co-led the benchmark data organization, implemented most model query and evaluation pipelines under a consistent and unified framework, and conducted the experiments and analyses. Tianhao Liang also maintained the evaluation pipeline and implemented the code execution metric. Tianhao Liang made significant efforts in data quality control, error analysis, and creating figures and tables for the paper.
Sherman Siu made contributions to the benchmark construction process, including task reviewing, annotator coordination, data quality control, metric implementation, and code maintenance. Sherman Siu contributed and designed a bunch of complex and novel planning tasks. Sherman Siu also analyzed the number of examples per task to investigate the variance of the benchmark score and contributed to writing the main paper.
Zhengqing Wang contributed around 40 tasks, including several complex traditional computer vision tasks. Zhengqing Wang organized the benchmark statistics for the Appendix, participated in error case analysis, and developed the project page.